OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

OpenStudy (waleed_imtiaz):

u know the formula to calculate the angle ???

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

do you?

OpenStudy (waleed_imtiaz):

the formula is : cos(theta) = u.v/|u| |v| In the numerator its the dot product of u and v and in the denominator its the magnitude of u and v seperately...and then take cos inverse to find theta...

OpenStudy (anonymous):

okay so numerator = 36?

OpenStudy (anonymous):

denominator = 12.92

OpenStudy (waleed_imtiaz):

numerator is right.

OpenStudy (waleed_imtiaz):

8.99*12.72 multiply these two u will get the denominator ..got it?

OpenStudy (anonymous):

im confused with the denominator

OpenStudy (anonymous):

oh okay so the denominator is 114.3528

OpenStudy (waleed_imtiaz):

okey then first calculate magnitude of u ... |u|= \[\sqrt{8^{2}+4^{2}}\] u will get 8.99

OpenStudy (waleed_imtiaz):

yes now u are rite

OpenStudy (anonymous):

so the angle is 71.65?

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (waleed_imtiaz):

I can try...

OpenStudy (anonymous):

Find the fourth roots of 256(cos 240° + i sin 240°).

OpenStudy (waleed_imtiaz):

first simplify the expression ..and then do like this....if i have to find the fourth root of unity then I will do x^4=1 x^4-1=0 simplify it and u will get x=1,-1, iota, -iota.... in the same way first write the value of cos 240 and sin 240 and make it simple to operate and then do like above as I told

OpenStudy (anonymous):

I don't understand that at all.. lol